Hi
I am sorry, I am truly sorry, I only resubmitting this because this is my first time, and I am not sure I did it right.
Soybean meal is 16%protein, cornmeal is 8%. How many pounds of each is needed to make 320 pounds that is 13% protein?
I got x+y=320
0.08x+0.16y=320() (0.13)
since no variable appears alone I used the elimination method.
8y=2547
318.375=320
y=318.3
x=1.625
Something just doesn't seem right about my answer. Please help, and thank you. Math is my WaterLoo and I thought I was working this correctly but the answers never seem to make any sense.

Answer by mananth(16949)

Soybean meal is 16%protein, let quantity be x pounds
cornmeal is 8%. will be (320-x) pounds
TOTAL 320 pounds that is 13% protein?
..
0.16x+0.08(320-x)= 0.13*320
0.16x+25.6 -0.08x=41.60
0.08x=41.6-25.6
0.08x=16
x= 200 pounds of soybean
Balance will be corn 320-200=120 pounds